OpenStudy (anonymous):

x-1= the square root of 3x+7

OpenStudy (colouringme):

the x in this equation has ranges, \[x \ge1\] and \[x \ge-\frac{ 7 }{ 3 }\] becauset there is no negative value of results from square root - ing something 1. square both side \[x - 1 = \sqrt{3x + 7}\] \[(x - 1)^{2} = (\sqrt{3x + 7})^{2}\] \[x ^{2} - 2x + 1 = 3x + 7\] \[x ^{2} - 5x - 6 = 0\] 2. factorise the new equation \[(x - 6) (x + 1)=0\] 3. there are two possible answers : x = 6 or x = -1 but x = 1 wouldn't fit the equation. so, x = 6 Sorry if I was wrong, but the answers fit the equation :)